A Manufacturer produces two products, the Klunk and the Klick. Klunk has a contribution



to profit of $3, and the Klick $4 per unit. The manufacturer wishes to establish the weekly



production plan that maximizes profit. Production of these products is limited to machine,



labor and material constraints. Each Klunk requires four hours machining, four hours labor



and one kilogram of material, where as each Klick requires two hours machining, six hours



labor and one kilogram of material. Machining and labor has a maximum of one hundred



and one hundred and eighty hours available, and total material available is forty kilograms.



Because of a trade agreement, sales of Klunk are limited to a weekly maximum of twenty



units and to honor an agreement with an old established customer at least ten units of Klick



must be sold each week.



i. Determine graphically using linear programming a suitable production mix of Klunk



and Klick. [12]



ii. What will be the company’s maximum profit?



1
Expert's answer
2021-12-13T14:16:17-0500

Let "x=" the number of units of Klunk to be produced, "y ="the number of units of Klick to be produced.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & Machine & Labour & Material \\ 
 & hours & hours & (kg) \\ 
\hline
 Klunk & 4x & 4x & x \\ 
 Klick & 2y & 6y & y \\ 
 \hdashline
Total & & & \\
 available & 100 & 180 & 40 \\
 per\ week & & & \\
\end{array}"



Maximize "z=3x+4y"

Subject to the constraints


"4x+2y\leq 100"

"4x+6y\leq180"

"x+y\leq40"

"0\leq x\leq20"

"y\geq 10"


i. The solution set of this system is the shaded region in the diagram


"4x+2y= 100=>y=-2x+50""x=20, y=10"

Point "C(20, 10)"



"4x+6y= 180=>y=-\dfrac{2}{3}x+30""x=0, y=30"

Point "A(0, 30)"



"x=0, y=10"

Point "D(0, 10)"



"-2x+50=-\dfrac{2}{3}x+30""\dfrac{4}{3}x=20""x=15, y=20"

Point "B(15, 20)"


ii.


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 Vertices & z=3x+4y \\ \hline
 A(0,30) & z=3(0)+4(30)=120 \\
 \hdashline
 B(15,20) & z=3(15)+4(20)=125 \\
 \hdashline
 C(20,10) & z=3(20)+4(10)=100 \\
 \hdashline
 D(0,10) & z=3(0)+4(10)=40 \\
 \hdashline

\end{array}"



Number of units of Kluck is 15 and number of units of Klick is 20.

Maximum profit is "z(x=15, y=20)=\$125."